OpenStudy (anonymous):

(Logarithms) solve for x : \[\x^{3/4(log_{2}x)^{2}+log2(x)−5/4}=sqrt{2}\]

OpenStudy (anonymous):

hmm this is complicated.

OpenStudy (anonymous):

\[x^{3/4(\log_{2}x)^{2}+\log_{2}(x) - 5/4} = \sqrt{2}\]

OpenStudy (anonymous):

oh my....I wasn't kidding...

OpenStudy (anonymous):

I got the equation - \[3t^{3} +4t^{2}-5t-2=0\] where, \[t=\log_{2}x\]

OpenStudy (anonymous):

i'm confused what is the actual question... where is the starting piont

OpenStudy (anonymous):

2nd post.

OpenStudy (anonymous):

is that x to power of everything ther

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

If you have x in the base and in the exponents, then the best you can possibly hope for is an approximation since an exact answer is pretty much impossible.

OpenStudy (anonymous):

that means none of you knows the answer.